OpenStudy (anonymous):

say you have 3x/3x+1. can you cancel the 3x?

OpenStudy (anonymous):

yes i think

OpenStudy (anonymous):

\[\frac{ 3x }{ 3x+1 }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

your question is like this or not

OpenStudy (anonymous):

its gona be 1/1 then =1

OpenStudy (anonymous):

tadaaa

OpenStudy (anonymous):

thanks so much! i'm solving a problem using implicit differentiation, and i needed to simplify my answer. I couldn't remember all of the canceling rules though. Thanks for your help!

hartnn (hartnn):

@gabby8292 NO YOU CANNOT CANCEL 3x

OpenStudy (anonymous):

if my answer was y'=[ ysin(xy)]/[xsin(xy)+ (.5)squareroot y'] I can't cancel the sin(xy) correct?

hartnn (hartnn):

no, you cannot. thats the simplest form.

hartnn (hartnn):

\(\huge \frac{a}{a+b}\ne \frac{1}{b}\)

OpenStudy (anonymous):

thanks for your help!

hartnn (hartnn):

welcome ^_^ and \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\)